Shift Summary for 01/05/2000 DAY shift


Subject: Shift Summary for 01/05/2000 DAY shift
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:02:13 CST


              Shift Summary for day shift, 01/05/00

On Shift: ER, TA

Hours of beam available at all intensities :2.0h
Hours of beam available at nominal intensity :2.0h
Hours of E832 data (special runs not included):0.8h
Hours of special runs (muon run) :0.2h

Accounting of lost time:
fresh start, replace a memory module in plane 4

========================================================================
Runs ended this shift:
 Run 15401 - BEAM_MU_4E12
 Run 15402 - 0.6M events - BEAM_799
 Run 15403 - 0.1M events - BEAM_799

========================================================================
Tapes brought to FCC this shift: None

========================================================================
Accesses during shift: None

========================================================================
Overview of shift:
 ran muon-run, got beam for a short time but accelerator spill signal
 prevented us for continuing collecting data

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15401 11:20 start muon run

              12:15 new sdaq version was made (v6_06)

15402 13:05 restore BA HV. start BEAM_799 run w/ neg AN4
              13:45 trigger stopped; problem with DDD 4 busy;
                        Jason L. attempts to debug it.

15403 14:00 fresh start new run
              14:05 same DDD 4 problem. JL replaces stream 6, plane 4
                        3rd memory (128MB) module.

15404 14:15 fresh start new run
              14:35 Beam is gone; No spill signal, seems to be a problem
                        with timeline generator; no time estimate

========================================================================
Follow up Notes on Previous Shift

 o standard muon-run ---> done
 o restore BA HV from DBase ---> done
 o standard E799 runnning NM4AN Neg ---> tried to do it...

========================================================================
Items to Follow up on Next Shift

 o take data, if possible...
   

 



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:02:23 CST